OpenStudy (anonymous):

explain the steps that are necessary to solve 4x - 2(x + 3) = x + 1

OpenStudy (anonymous):

First you must distribute the -2. Do you know what the distributive property is?

OpenStudy (anonymous):

Yeah, but how about you tell me!

OpenStudy (anonymous):

The distributive property is when you take the number on the outside of the parenthesis and multiply it with the numbers inside the parenthesis. So after using the distributive property your equation will look like this : 4x - 2x - 6 = x + 1

OpenStudy (anonymous):

sooo whats the answer?

OpenStudy (anonymous):

From here you combine like terms. 2x - 6 = x + 1 Subtract x from both sides of the equation. x - 6 = 1 Add 6 to both sides of the equation. x = 7 So your solution is x = 7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your welcome.